From mar@roe.ac.uk Fri Apr 27 14:13:04 2012 Date: Fri, 27 Apr 2012 13:53:57 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Andrew Lawrence , Bob Mann , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Norman Gray , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives project team meeting minutes, 27/4/2012 Present: MAR, MSH, RPB, RGM, ETWS, NG Apologies: NCH, JPE, AL, KTN, STV, NJC, RSC DoNM: 10am Friday May 11th 2012 in the VH (NCH in the chair) Actions discharged this time: ----------------------------- ACTION: MAR to update VSA pages with details of public survey releases. Done for all current public releases. ACTION: NCH to contact VIKING PI to clarify what needs doing by when, noting that we may struggle a bit on the list-driven aspects NCH had sent an email to Will raising this and suggesting a telecon. No reply to date. ACTION: NJC to contact VMC PI stating WFAU's position on dereddening and variability classification in ESO phase III deliveries Maria-Rosa replied to NJC's missive basically agreeing with our sentiments. With regard to variable classification they do have catalogues of Cepheids and RR Lyrae stars in the two completed VMC tiles so these might be submitted to ESO pending further discussions. ACTION: NCH to contact VVV PIs to discuss requirements for the phase III deliverables Dante replied agreeing with our pragmatic approach of suggesting more realistic targets and suggested drafting a response in May after the OPC has met. ACTION: NJC to contact VIDEO PI to clarify what (if anything) needs to be done for their phase III deliverables. No reply from Matt so the assumption is they have it in hand (or the reply just went to NJC). ACTION: RPB to put UKIDSS DR8 world-public, sans GPS and UDS. Done (only GPS excluded as UDS is past its proprietary period too) ACTION: RPB to activate GPS access in world-public UKIDSS DR7 Done. ACTION: NJC to contact UHS survey scientist Simon Dye to open up discussions as to requirements for UHS curation/dissemination. Simon has supplied some details and will be back in contact following discussions with the UHS working group. Actions partly discharged but continuing: ----------------------------------------- ACTION: RPB to roll out latest curation client / DB server share configuration for routine ops use. Continues; ETWS working on CU4 code. Actions carried forward from 13/04/12 meeting --------------------------------------------- ACTION: RPB to set up WSA mirror/sync for flat file DBs using MS SQL replication. Continues. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. ACTION: RPB to complete the LTO-5 tape backups of VVV and SSA DBs and SSS flat file datasets. Continues; RPB has sourced the required SAS interface card and an order is pending. Specific points and new actions: -------------------------------- Project management: WFAU's WSA report has been submitted to the UKIRT board. Prior to the grant renewal submission at the end of June we would have the opportunity to impress two visitors. A Korean delegate attending the UKIRT board meeting would spend a day with WFAU, and Richard Hook (ESO) will visit on the 21st May to discuss publicising VISTA. WFCAM, VISTA and VST updates: Nothing to report this time Comments and issues arising from CASU minutes: No new minutes this time. Networking: Currently experiencing slow transfer speeds, under investigation. WSA/VSA/OSA/GES Operations: ETWS reported: -- Transferred January 2012 VISTA data and started transfer of the February VISTA data. Transfer speeds are very slow, only a quarter of the speed we get normally, and MSH is looking into it. RPB noted that Ramses12 is now the VVV load server and that the VVV release is being cu19'ed. The VIKING ESO release is also still running. Hardware and Systems: Apart from the problem with disk14 earlier in the week everything else is running fine. MSH noted that we would need a new InfiBand switch shortly. As these are over 10K it would be better to buy this sooner rather than later. Survey Data Release: UKIDSS releases as noted in the discharged actions above. Software: ETWS reported: -- Working on the ingester scripts to include the changes of the data shares. -- Installing software on the new servers neferefre and ini. Nothing of note this time. Non-survey Data Release: RPB reported the release of 10 non-survey databases. Astrogrid deployment & Data Analysis services: MSH would be attending the VO-ICE and IVOA meetings in May. Miscellaneous: NG reported progress on generating UCDs from locating where table attributes lie in a model of the observations. RGM reported that we were hoping to host an expert in MonetDB for a 3 month period starting in June. -- Scanned by iCritical.